OpenStudy (anonymous):

Let (9,10) be a point on the terminal side of theta. Find cos theta, sec theta, and cot theta.

OpenStudy (anonymous):

http://www.mathopenref.com/trigterminalside.html Ok so you have a triangle with 9 being the x component and 10 being the y component. Pythagorean Theorem yields sqrt(9^2 + 10^2) = 13.454 so the hypotenuse is 13.454 units long. cos theta = adj/hyp = 9/13.454 sec theta = hyp/adj = 13.454/9 cot theta = adj/opp = 9/10 You can simplify if you like